HAROUN COMPANY
Comparative Income Statements
For Years Ended December 31, 2019–2013
($ thousands) 2019 2018 2017 2016 2015 2014 2013
Sales $ 2,778 $ 2,433 $ 2,214 $ 2,029 $ 1,893 $ 1,760 $ 1,443
Cost of goods sold 1,997 1,625 1,398 1,223 1,137 1,063 847
Gross profit 781 808 816 806 756 697 596
Operating expenses 595 465 427 315 273 269 224
Net income $ 186 $ 343 $ 389 $ 491 $ 483 $ 428 $ 372
HAROUN COMPANY
Comparative Balance Sheets
December 31, 2019–2013
($ thousands) 2019 2018 2017 2016 2015 2014 2013
Assets
Cash $ 160 $ 211 $ 219 $ 224 $ 232 $ 229 $ 237
Accounts receivable, net 1,146 1,204 1,090 836 736 697 492
Merchandise inventory 4,147 3,019 2,638 2,223 1,997 1,696 1,230
Other current assets 106 96 59 106 89 91 47
Long-term investments 0 0 0 327 327 327 327
Plant assets, net 5,075 5,055 4,422 2,493 2,579 2,292 1,967
Total assets $ 10,634 $ 9,585 $ 8,428 $ 6,209 $ 5,960 $ 5,332 $ 4,300
Liabilities and Equity
Current liabilities $ 2,673 $ 2,247 $ 1,475 $ 1,227 $ 1,064 $ 1,007 $ 649
Long-term liabilities 2,857 2,488 2,421 1,124 1,149 1,244 933
Common stock 1,935 1,935 1,935 1,720 1,720 1,505 1,505
Other paid-in capital 484 484 484 430 430 376 376
Retained earnings 2,685 2,431 2,113 1,708 1,597 1,200 837
Total liabilities and equity $ 10,634 $ 9,585 $ 8,428 $ 6,209 $ 5,960 $ 5,332 $ 4,300


Required:
1. Complete the below table to calculate the trend percents for all components of both statements using 2013 as the base year.
2. Refer to the results from part 1. (a) Did sales grow steadily over this period? (b) Did net income as a percent of sales grow over the past four years? (c) Did inventory increase over this period?
Complete the below table to calculate the trend percents for all components of comparative income statements using 2013 as the base year. (Round your percentage answers to 1 decimal place.)

HAROUN COMPANY
Income Statement Trends
For Years Ended December 31, 2019–2013
2019 2018 2017 2016 2015 2014 2013
Sales % % % % % % 100.0 %
Cost of goods sold 100.0
Gross profit 100.0
Operating expenses 100.0
Net income % % % % % % 100.0 %
Answer #1
2019 2018 2017 2016 2015 2014 2013
Sales 193% 169% 153% 141% 131% 122% 100%
Cost of goods sold 236% 192% 165% 144% 134% 126% 100%
Gross Profit 131% 136% 137% 135% 127% 117% 100%
Operating Expenses 266% 208% 191% 141% 122% 120% 100%
Net Income 50% 92% 105% 132% 130% 115% 100%

Here is the formula version

C D E F G H I 2019 =C2-1 =D2-1 =E2-1 =F2-1 =G2-1 =H2-1 1625 Sales 2778 COGS 1997 GP =C4-C5 Operating Expenses 595 Net Income

2.

a) Yes, there is a consistent growth in sales over the period. The growth was 22% in 2014, then there was further growth in the following years to the point where it grew to almost double in 2019.

b)

NI as a percent of sales 7% 14% 18% 24% 26% 24% 26%

No, net income as a percent of sales did not show growth. It was on a decline.

c)

Yes, as can be seen from the Balance Sheet, ending inventory consistently increased over the period in question.